OpenStudy (anonymous):

The perimeter of a regular hexagon is 108. What is its area?

OpenStudy (anonymous):

108/6=18, length of hexagon edge A=(3*sqrt(3)/2)*t^2, where t is the edge length A=841.7767

OpenStudy (anonymous):

omar_86 i cannot fnd that answer in my list.

OpenStudy (anonymous):

what options u have?

OpenStudy (anonymous):

30, 41,31,0

OpenStudy (anonymous):

and the perimeter is 108?

OpenStudy (anonymous):

sorry i gave the choices to the wrong question.

OpenStudy (anonymous):

486*square root 3, 81*square root 3, 729, 729*square root 3

OpenStudy (anonymous):

it's 486*square root 3

OpenStudy (anonymous):

can you please explain for clarification thanks.

OpenStudy (anonymous):

the law of hexagon area is: 3*sqrt(3) -------- * t^2 2 where t is the edge length, ok till now?

OpenStudy (anonymous):

thank u im fine

OpenStudy (anonymous):

we divide the whole perimeter length by 6, since hexagon, to get the edge length, in this case 108/6=18

OpenStudy (anonymous):

then substitute 18 into the previous equation instead of t